# Semantic-Aware Adaptive Video Streaming Using Latent Diffusion Models for Wireless Networks

### v2

@@ IV-A Channel-Aware Video Streaming Bitrate Selector
  • Neural Network: For each episode $e$, the ABS learning agent determines a suitable bitrate for the next chunk via a NN. The input features of this NN in $k$-th chunk are ($S_{k}=\{C_{k},M_{k},F_{k}\}$) and are detailed as follows: 1) Past Wireless Channel Estimate: The learning agent uses the past $t$ chunks’ channel status vector $C_{k}=\{c_{k-t},\dots,c_{k}\}$ as input to the NN, where $c_{i}$ represents the channel throughput for video chunk $i$. 2) Latent Content: To detect the diversity of video content, the learning agent uses $M_{k}=\{N_{k+1},V_{k+1}\}$, where $N_{k+1}$ and $V_{k+1}$ denote the size of each bitrate of the next chunk $(k+1)$ and the perceptual quality metrics (semantic ambiguities) for each bitrate of the next chunk, respectively. 3) Player Rollout Playback: Rollout information is collected by the player on the edge user’s side and is defined as $F_{k}=\{v_{k-1},B_{k},D_{k},m_{k}\}$. Here, $v_{k-1}$ represents the video quality of the last video chunk selected, while $B_{k}$, $D_{k}$, and $m_{k}$ represent the buffer utilization, loading time, and the normalized remaining chunks, respectively, for the past $t$ chunks.
@@ IV-B LDM-enabled Compression, Denoising, and Recontruction
For the remainder of LD-ABS, to achieve higher compression of streaming video, we consider using LDM to compress the original reference frames into latent reference I frames. Meanwhile, we only retain the adjustment metadata for the predictive B and P frames. On the user side, latent reference frames are denoised to restore reference frames, and predictive frames are restored using conditional LDM and VFI. Recognizing the challenges faced by LDM-enabled real-time adaptive multi-band streaming, we propose a collaborative end-to-end LDM-enabled streaming framework. This framework aims to deliver high-resolution, low-latency streaming video while conserving communication bandwidth and storage with adaptive time-varying appropriate bitrate.
Specifically, the proposed framework involves the following key components:
@@
Figure 3: LDM-enabled adapted bitrate video streaming framework (LDM-ABS)
@@ IV-C Key Benefits of LD-ABS
Overall, Fig. 3 depicts the structure of the proposed LDM-ABS framework , where $\bm{z}$ is the semantic latent features of video frames $\bm{x}$ encoded by VAE, $T$ represents the total number of forward steps of DM, $t$ denotes the $t$-th step, $n$ represents the $n$-th frame of the video stream with total $N$ frames, $\bm{\theta}$ is the parameters of the DM’s denoising U-Net neural network [4], which include 128 base channels in the base layer of the encoder with channel multipliers [1, 2, 4, 8], $r$ represents the starting point for denoising received key frame semantic latent feature signals in the wireless channel, which is related to the channel state, $s$ represents the starting step for fine-tuning non-key frame semantic latent feature, and $\bm{c}$ is the condition for the conditional LDM, i.e., the small-size meta-data being transmitted. Furthermore, $q(\cdot|\cdot)$ denotes the forward process of the DMs, and $p_{\bm{\theta}}(\cdot|\cdot)$ represents the reverse denoising process.
@@ V-C QoE Performance Comparison
As depicted in Fig. 5, LD-ABS outperforms the other benchmark with higher average QoE improvements across different traces. The other benchmark frameworks show comparable loss compared with our proposed LD-ABS. With CSI error, LD‑ABS loses <3 % QoE, four times less than the closest competitor. The reason is two‑fold: (i) denoising is performed in a low‑dimensional latent space with only a handful of reverse steps, so the error does not propagate deeply; (ii) the subsequent conditional diffusion for B/P‑frames is anchored to the already‑refined I‑frame, effectively absorbing residual noise. Fig. 6 illustrates a sample streaming simulation within a chunk. This chunk consists of one I-frame followed by a P-frame, another P-frame, and a B-frame. Instead of transmitting the entire frames through a noisy wireless channel, only lightweight adjustment metadata is transmitted. The LDM then reconstructs the respective latent frames based on the recovered latent I frame ($z_{o}^{n}$) and the received adjustment metadata, generating the corresponding adjusted frames accordingly.
Figure 6: Sample video chunk processing using the proposed LDM-ABS framework.

### v3

@@ IV-A Channel-Aware Video Streaming Bitrate Selector
[…]
@@ IV-B LDM-enabled Compression, Denoising, and Recontruction
For the remainder of LD-ABS, to achieve higher compression of streaming video, we consider using LDM to compress the original reference frames into latent reference I frames. Meanwhile, we only retain the adjustment metadata for the predictive B and P frames. On the user side, latent reference frames are denoised to restore reference frames, and predictive frames are restored using conditional LDM and VFI. Recognizing the challenges faced by LDM-enabled real-time adaptive multi-band streaming, we propose a collaborative end-to-end LDM-enabled streaming framework. This framework aims to deliver high-resolution, low-latency streaming video while conserving communication bandwidth and storage with adaptive time-varying appropriate bitrate. Overall, Fig. 3 depicts the structure of the proposed LD-ABS framework , where $\bm{z}$ is the semantic latent features of video frames $\bm{x}$ encoded by VAE, $T$ represents the total number of forward steps of DM, $t$ denotes the $t$-th step, $n$ represents the $n$-th frame of the video stream with total $N$ frames, $\bm{\theta}$ is the parameters of the DM’s denoising U-Net neural network [4], which include 128 base channels in the base layer of the encoder with channel multipliers [1, 2, 4, 8], $r$ represents the starting point for denoising received key frame semantic latent feature signals in the wireless channel, which is related to the channel state, $s$ represents the starting step for fine-tuning non-key frame semantic latent feature, and $\bm{c}$ is the condition for the conditional LDM, i.e., the small-size meta-data being transmitted. Furthermore, $q(\cdot|\cdot)$ denotes the forward process of the DMs, and $p_{\bm{\theta}}(\cdot|\cdot)$ represents the reverse denoising process. Specifically, the proposed framework involves the following key components:
@@
Figure 3: LDM-enabled adapted bitrate video streaming framework (LD-ABS)
@@ V-C QoE Performance Comparison
As depicted in Fig. 5, LD-ABS outperforms the other benchmark with higher average QoE improvements across different traces. The other benchmark frameworks show comparable loss compared with our proposed LD-ABS. With CSI error, LD-ABS loses <3 % QoE, four times less than the closest competitor. The reason is two-fold: (i) denoising is performed in a low-dimensional latent space with only a handful of reverse steps, so the error does not propagate deeply; (ii) the subsequent conditional diffusion for B/P-frames is anchored to the already-refined I-frame, effectively absorbing residual noise. Fig. 6 illustrates a sample streaming simulation within a chunk. This chunk consists of one I-frame followed by a P-frame, another P-frame, and a B-frame. Instead of transmitting the entire frames through a noisy wireless channel, only lightweight adjustment metadata is transmitted. The LDM then reconstructs the respective latent frames based on the recovered latent I frame ($z_{o}^{n}$) and the received adjustment metadata, generating the corresponding adjusted frames accordingly.
Figure 6: Sample video chunk processing using the proposed LD-ABS framework.